News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 31 March 2017 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

Real-time procedures were executed on Mar 27 in preparation for a 20-minute-long, penumbral Lunar eclipse on Mar 28: one to dump and clear the EPS glitch counters, one to activate SCS 29, and one to set the short eclipse flag to 'true'. Chandra passed through the eclipse with nominal power and thermal performance. Real-time procedures were executed following the eclipse on Mar 28 to disable SCS 29, to set the short eclipse flag to 'false', and to dump and clear the EPS glitch counters.

A Chandra press release was issued on Mar 30 describing observations of an X-ray transient located in the Chandra Deep Field-South. The source was seen to brighten by at least a factor of a thousand in a few hours. Possible explanations include a gamma ray burst seen in an unusual way or a medium-sized black hole shredding a white dwarf star. For details see:

The schedule of targets for the next week is shown below and includes an observation of HD80404 coordinated with HST, and observations of SgrA* coordinated with the Event Horizon Telescope.

        Radiation Belts
        Off-field2             ACIS-I       Apr  3
        PROCYON                HRC-I
        W0615-5716             ACIS-S
        TWA33                  ACIS-S
        A1795                  ACIS-I
        HZ43                   HRC-S /LETG  Apr  4
        J1207-3900             ACIS-S
        IGRJ17315-3221         ACIS-I
        Radiation Belts                     Apr  5
        SDSSJ125415.55+48085   ACIS-S
        betaLeo                HRC-I
        A1795                  ACIS-I
        SgrA*                  ACIS-S       Apr  6
        WD1145+017             ACIS-S
        NGC4450                ACIS-S
        J1111-2655             ACIS-S
        SgrA*                  ACIS-S       Apr  7
        W2210-3507             ACIS-S
        2MASSJ20125995+0112    ACIS-S
        Radiation Belts                     Apr  8
        Abell1795outskirts     ACIS-I
        A1795                  ACIS-I
        CygnusA-WHotspot       ACIS-I
        A1795                  ACIS-I       Apr  9


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports